OpenStudy (anonymous):

7 multiplied by 3 multiplied by 2 whole over 7 multiplied by 5, the whole raised to the power of 2 multiplied by 7 to the power of 0 over 5 to the power of negative 3, whole to the power of 3 multiplied by 5 to the power of negative 9? do you know the equation

OpenStudy (anonymous):

@math&ing001

OpenStudy (johnweldon1993):

\[\large (\frac{7\times 3 \times 2}{7\times 5})^2 \times (\frac{7^0}{5^{-3}})^3 \times 5^{-9}\] Like that?

OpenStudy (anonymous):

lol I am helping a friend I didnt understand the by 2

OpenStudy (anonymous):

ok it is right thanks so much . he Approves

OpenStudy (welshfella):

calculate what inside the parentheses first

OpenStudy (johnweldon1993):

Oh lol no problem!